# From Sparse Dependence to Sparse Attention: Unveiling How Chain-of-Thought Enhances Transformer Sample Efficiency

### v1

@@ A.4.1 Auxiliary Statistics
$\gamma_{i,b,i^{\prime},b_{2}}$ $\frac{1}{i^{\prime}}\left(\sum_{r=1}^{2m}\bm{1}\left(\nu_{r,i,b}>0\right)\bm{1}\left(\nu_{r,i^{\prime},b_{2}}>0\right)\right)-\frac{1}{i^{\prime}}\frac{1+\bm{1}\left(i=i^{\prime},\ b=b_{2}\right)}{2}m$ $\mathcal{O}\left(\frac{\sqrt{m\log\left(\tfrac{n}{\delta}\right)}}{n}\right)$
@@
For each batch, the data inside the batch is approximately balanced. With probability at least $101-\delta/$ 10, the following bounds hold for all $t\in[3]$, $i,j\in[n+k]$, $b,b_{1},b_{2}\in\{0,1\}$, and $r\in[m]$:
@@
A.4.2 First Step: Configuring the MLPs
@@
$\left|\left\langle W^{0}_{r,1:d},\mathbf{E}(\textbf{b})[i]\right\rangle-\nu_{r,i,\textbf{b}[i]}\right|\leq\sqrt{\frac{2\log(\frac{100Bm}{\delta})k}{d}}\epsilon$
By Lemma 5.1, $\sqrt{\frac{2k}{d}\log(Bm/100\delta)}<0.01$, we can conclude that
@@
$\frac{d\mathcal{L}^{(0)}}{dW_{r,1:d}}=$ $\frac{1}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(t,s)}[i+1]}\nabla_{W_{r,1:d}}\mathcal{T}[w^{(t)}](\textbf{b}^{(t,s)})[i]$
$=$ $\frac{h_{r}}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(t,s)}[i+1]}e_{i,\textbf{b}^{(t,s)}[i]}\bm{1}(\nu_{r,i,\textbf{b}^{(t,s)}[i]}>0).$
@@
$\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(t,s)}[i+1]}e_{i,\textbf{b}^{(t,s)}[i]}\bm{1}(\nu_{r,i,\textbf{b}^{(t,s)}[i]}>0)$
$=$ $\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}\sum_{b=0}^{1}(-1)^{\textbf{b}^{(t,s)}[i+1]}\bm{1}(\textbf{b}^{(t,s)}[i]=b)e_{i,b}\bm{1}(\nu_{r,i,b}>0)$
$=$ $\sum_{i=n+1}^{n+k}\sum_{b=0}^{1}e_{i,b}\bm{1}(\nu_{r,i,b}>0)\sum_{s=1}^{B}(-1)^{\textbf{b}^{(t,s)}[i+1]}\bm{1}(\textbf{b}^{(t,s)}[i]=b)$
@@
$\left|\left\langle\frac{h_{r}}{B}\sum_{i=n+1}^{n+k}\sum_{b=0}^{1}\bm{1}(\nu_{r,i,b}>0)\delta_{0,i,b}e_{i,b},e_{i,b}\right\rangle\right|\leq\frac{1}{2mB}\sup\delta_{t,i,b}\left(1+\sqrt{\frac{2k\log\left(\frac{200m}{\delta}\right)}{d}}\right)$
@@
For $\bar{\alpha}$ and $\psi$ defined in Section A.4.1, When $\mathcal{E}_{1}$ defined in Lemma 11 happens, for all $i^{\prime}\in[n+1,n+k+1],b^{\prime}\in\{0,1\}$, the gradient on $W_{r,d+1:2d}$ satisfies that,
@@
$\frac{d\mathcal{L}^{(0)}}{dW_{r,d+1:2d}}=$ $\frac{1}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(t,s)}[i+1]}\nabla_{W_{r,1:d}}\mathcal{T}[w^{(t)}](\textbf{b}^{(t,s)})[i]$
$=$ $\frac{h_{r}}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(t,s)}[i+1]}\mathcal{A}[A^{(t)}](\mathbf{E}(\textbf{b}^{(t,s)}))[i]\bm{1}(\nu_{r,i,\textbf{b}^{(t,s)}[i]}>0)$
$=$ $\frac{h_{r}}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(t,s)}[i+1]}\bm{1}(\nu_{r,i,\textbf{b}^{(t,s)}[i]}>0)\frac{1}{i}\sum_{j=1}^{i}\sum_{b_{1}=0}^{1}\bm{1}\left(\textbf{b}[j]=b\right)e_{j,b_{1}}$
@@
$+\frac{\sqrt{2\log(\frac{400mn}{\delta})}}{2mB\sqrt{d}}\left(\sqrt{\sum_{i=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i}\bm{1}(\nu_{r,i,b_{2}}>0)\bar{\alpha}_{i,S[i],b_{1},b_{2}}^{2}}+\sqrt{\sum_{j=1}^{n+k}\sum_{b_{1}=0}^{1}\psi_{0,r,j,b_{1}}^{2}}\right)$
@@
$\frac{\sqrt{2\log(\frac{400mn}{\delta})}}{2mB\sqrt{d}}\sqrt{\sum_{i=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i}\bm{1}(\nu_{r,i,b_{2}}>0)\bar{\alpha}_{i,S[i],b_{1},b_{2}}^{2}}\leq\frac{\sqrt{2k\log(\frac{400mn}{\delta})}}{2mn\sqrt{d}}\leq\frac{1}{200mn},$
@@
A.4.3 Step 2: Configuring the Attention
@@
For $\alpha,\beta,\gamma,\zeta$ defined in Section A.4.1, when the event in Lemma 17 happens, for any $i$ and $b$,
$\kappa_{1,i,b}=$ $\frac{\eta_{0}}{8Bmi}\sum_{b_{1}=0}^{1}\bar{\alpha}_{i,S[i],b_{1},b}e_{S[i],b_{1}}+\frac{\eta_{0}}{4Bm^{2}}\left(\sum_{i^{\prime}=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\gamma_{i,b,i^{\prime},b_{2}}\bar{\alpha}_{i^{\prime},S[i^{\prime}],b_{1},b_{2}}e_{S[i],b_{1}}\right)$
@@
$\frac{d\mathcal{L}^{(0)}}{dW_{r,d+1:2d}}=\frac{h_{r}}{B}\sum_{i=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i}\bm{1}(\nu_{r,i,b_{2}}>0)\bar{\alpha}_{i,S[i],b_{1},b_{2}}e_{S[i^{\prime}],b_{1}}$ $+\frac{h_{r}}{B}\sum_{j=1}^{n+k}\sum_{b_{1}=0}^{1}\psi_{0,r,j,b_{1}}e_{j,b_{1}}.$
@@
$\kappa_{1,i,b}=$ $\frac{\eta_{0}}{4Bm^{2}}\left(\sum_{i^{\prime}=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\left(\sum_{r=1}^{2m}\bm{1}\left(\nu_{r,i,b}>0\right)\bm{1}(\nu_{r,i^{\prime},b_{2}}>0)\right)\frac{1}{i^{\prime}}\bar{\alpha}_{i^{\prime},S[i^{\prime}],b_{1},b_{2}}e_{S[i],b_{1}}\right)$
@@
$+\frac{\eta_{0}}{4Bm^{2}}\left(\sum_{i^{\prime}=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\gamma_{i,b,i^{\prime},b_{2}}\bar{\alpha}_{i^{\prime},S[i^{\prime}],b_{1},b_{2}}e_{S[i],b_{1}}\right)$
@@
$\left|\frac{\eta_{0}}{8Bmi}\left\langle\sum_{b_{1}=0}^{1}\bar{\alpha}_{i,S[i],b_{1},b}e_{S[i],b_{1}}e_{j,b_{1}}\right\rangle-\frac{\eta_{0}}{8Bmi}\bar{\alpha}_{i,S[i],b_{1},b_{2}}\bm{1}(j=S[i])\right|\leq$ $\frac{2\sqrt{\log(200n/\delta)}}{8mn\sqrt{d}}.$
@@
$\left|\langle H_{i,b},e_{j,b_{1}}\rangle-\frac{\eta_{0}}{8Bmi}\bar{\alpha}_{i,S[i],b_{1},b_{2}}\bm{1}(j=S[i])\right|\leq\frac{10\eta_{0}}{mn}\frac{\sqrt{nk}}{\sqrt{Bd}}\log(300mnB/\delta).$
@@
$\frac{d\ell(\mathcal{T}(\textbf{b})[i],\textbf{b}[i+1])}{dA}=(-1)^{\textbf{b}[i+1]+1}\sum_{j=1}^{i}\langle\kappa_{1,i,\textbf{b}[i]},e_{j,\textbf{b}[j]}\rangle\frac{\partial\mathrm{softmax}\left(\mathbf{E}(\textbf{b})^{T}A\mathbf{E}(\textbf{b})\right)[j,i]}{\partial A}.$
with,
@@
$=\sum_{r=1}^{2m}h_{r}W^{(1)}_{r}\bm{1}\left(\nu_{r,i,\textbf{b}[i]}>0\right)$
@@
$\frac{d\ell(\mathcal{T}(\textbf{b})[i],\textbf{b}[i+1])}{d\mathcal{A}(\mathbf{E}(\textbf{b}))[i]}$ $=(-1)^{\textbf{b}[i+1]}\left(\sum_{r=1}^{2m}h_{r}W^{(1)}_{r}\bm{1}\left(\nu_{r,i,\textbf{b}[i]}>0\right)\right)$
@@
$=$ $(-1)^{\textbf{b}[i+1]}\sum_{j=1}^{i}\sum_{r=1}^{2m}h_{r}\bm{1}\left(\nu_{r,i,\textbf{b}[i]}>0\right)\langle W^{(1)}_{r},\mathbf{E}(\textbf{b})[j]\rangle\frac{\partial\mathrm{softmax}\left(\mathbf{E}(\textbf{b})^{T}A\mathbf{E}(\textbf{b})\right)[j,i]}{\partial A}$
@@
$-\frac{dL^{(1)}}{dA}=$ $\frac{1}{B}\sum_{i=1}^{n}\sum_{j=1}^{i}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i}\alpha_{1,i,j,b_{1},b_{2}}\Delta_{1,i,j,b_{1},b_{2}}e_{j,b_{1}}e_{i,b_{2}}^{T}$
$-\frac{1}{B}\sum_{i=1}^{n}\sum_{j=1}^{i}\sum_{p=1}^{i}\sum_{b_{1}=0}^{1}\sum_{b_{2}=1}^{1}\sum_{b_{3}=0}^{1}\frac{1}{i^{2}}\mu_{1,j,i,m,b_{1},b_{2},b_{3}}\Delta_{1,i,j,b_{1},b_{2}}e_{k,b_{3}}e_{i,b_{2}}^{T}.$
@@
    $U_{j,p}$ can ve bounded in the same way as $T_{j}$ and we have that,
@@
$\left|\mathcal{A}[A^{2}](\mathbf{E}(\textbf{b}))[i]-e_{S[i],\textbf{b}[S[i]]}\right|<1/n^{9}.$
Proof.
This is a direct consequence of Lemma 21 with $\frac{\eta_{0}\eta_{1}}{256mn^{2}}>10\log n$ (Lemma 5). ∎
A.4.4 Third Step: Moving MLP in Linear Regime

### v2

@@ B.4.1 Auxiliary Statistics
[…]
@@
For each batch, the data inside the batch is approximately balanced. With probability at least $101-\delta/$ 10, the following bounds hold for all $t\in[3]$, $i,j\in[n+k]$, $b,b_{1},b_{2}\in\{0,1\}$, and $r\in[m]$ for statistics defined in Table 1:
@@
B.4.2 First Step: Configuring the MLPs
@@
$\left|\left\langle W^{0}_{r,1:d},\mathbf{E}(\textbf{b})[i]\right\rangle-\nu_{r,i,\textbf{b}[i]}\right|\leq\sqrt{\frac{4\log(\frac{100Bm}{\delta})k}{d}}\epsilon$
By Lemma 5.2, $\sqrt{\frac{4k}{d}\log(100Bm/\delta)}<0.01$, we can conclude that
@@
$\frac{d\mathcal{L}^{(0)}}{dW_{r,1:d}}=$ $\frac{1}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(0,s)}[i+1]}\nabla_{W_{r,1:d}}\mathcal{T}[w^{(t)}](\textbf{b}^{(0,s)})[i]$
$=$ $\frac{h_{r}}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(0,s)}[i+1]}e_{i,\textbf{b}^{(t,s)}[i]}\bm{1}(\nu_{r,i,\textbf{b}^{(0,s)}[i]}>0).$
@@
$\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(0,s)}[i+1]}e_{i,\textbf{b}^{(0,s)}[i]}\bm{1}(\nu_{r,i,\textbf{b}^{(0,s)}[i]}>0)$
$=$ $\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}\sum_{b=0}^{1}(-1)^{\textbf{b}^{(0,s)}[i+1]}\bm{1}(\textbf{b}^{(0,s)}[i]=b)e_{i,b}\bm{1}(\nu_{r,i,b}>0)$
$=$ $\sum_{i=n+1}^{n+k}\sum_{b=0}^{1}e_{i,b}\bm{1}(\nu_{r,i,b}>0)\sum_{s=1}^{B}(-1)^{\textbf{b}^{(0,s)}[i+1]}\bm{1}(\textbf{b}^{(0,s)}[i]=b)$
@@
$\left|\left\langle\frac{h_{r}}{B}\sum_{i=n+1}^{n+k}\sum_{b=0}^{1}\bm{1}(\nu_{r,i,b}>0)\delta_{0,i,b}e_{i,b},e_{i,b}\right\rangle\right|\leq\frac{1}{2mB}\sup\delta_{0,i,b}\left(1+\sqrt{\frac{2k\log\left(\frac{200m}{\delta}\right)}{d}}\right)$
@@
For $\bar{\alpha}$ and $\psi$ defined in Section B.4.1, When $\mathcal{E}_{1}$ defined in Lemma 11 happens, for all $i^{\prime}\in[n+1,n+k+1],b^{\prime}\in\{0,1\}$, the gradient on $W_{r,d+1:2d}$ satisfies that,
@@
$\frac{d\mathcal{L}^{(0)}}{dW_{r,d+1:2d}}=$ $\frac{1}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(0,s)}[i+1]}\nabla_{W_{r,1:d}}\mathcal{T}[w^{(t)}](\textbf{b}^{(0,s)})[i]$
$=$ $\frac{h_{r}}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(0,s)}[i+1]}\mathcal{A}[A^{(t)}](\mathbf{E}(\textbf{b}^{(0,s)}))[i]\bm{1}(\nu_{r,i,\textbf{b}^{(0,s)}[i]}>0)$
$=$ $\frac{h_{r}}{B}\sum_{s=1}^{B}\sum_{i=n+1}^{n+k}(-1)^{\textbf{b}^{(0,s)}[i+1]}\bm{1}(\nu_{r,i,\textbf{b}^{(0,s)}[i]}>0)\frac{1}{i}\sum_{j=1}^{i}\sum_{b_{1}=0}^{1}\bm{1}\left(\textbf{b}[j]=b\right)e_{j,b_{1}}$
@@
$+\frac{\sqrt{2\log(\frac{400mn}{\delta})}}{2mB\sqrt{d}}\left(\sqrt{\sum_{i=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i^{2}}\bm{1}(\nu_{r,i,b_{2}}>0)\bar{\alpha}_{i,S[i],b_{1},b_{2}}^{2}}+\sqrt{\sum_{j=1}^{n+k}\sum_{b_{1}=0}^{1}\psi_{0,r,j,b_{1}}^{2}}\right)$
@@
$\frac{\sqrt{2\log(\frac{400mn}{\delta})}}{2mB\sqrt{d}}\sqrt{\sum_{i=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i^{2}}\bm{1}(\nu_{r,i,b_{2}}>0)\bar{\alpha}_{i,S[i],b_{1},b_{2}}^{2}}\leq\frac{\sqrt{2k\log(\frac{400mn}{\delta})}}{2mn\sqrt{d}}\leq\frac{1}{200mn},$
@@
B.4.3 Second Step: Configuring the Attention
@@
$\sup_{r}\left|\left\langle W^{1}_{r}-W^{0}_{r},[\mathbf{E}(\textbf{b})[i],\mathcal{A}[A^{1}](\mathbf{E}(\textbf{b}))[i]]\right\rangle\right|<3\epsilon/80.$
@@
For $\alpha,\beta,\gamma,\zeta$ defined in Section B.4.1, when the event in Lemma 17 happens, for any $i$ and $b$,
$\kappa_{1,i,b}=$ $\frac{\eta_{0}}{8Bmi}\sum_{b_{1}=0}^{1}\bar{\alpha}_{i,S[i],b_{1},b}e_{S[i],b_{1}}+\frac{\eta_{0}}{4Bm^{2}}\left(\sum_{i^{\prime}=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\gamma_{i,b,i^{\prime},b_{2}}\bar{\alpha}_{i^{\prime},S[i^{\prime}],b_{1},b_{2}}e_{S[i^{\prime}],b_{1}}\right)$
@@
$\frac{d\mathcal{L}^{(0)}}{dW_{r,d+1:2d}}=\frac{h_{r}}{B}\sum_{i^{\prime}=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i^{\prime}}\bm{1}(\nu_{r,i^{\prime},b_{2}}>0)\bar{\alpha}_{i^{\prime},S[i^{\prime}],b_{1},b_{2}}e_{S[i^{\prime}],b_{1}}$ $+\frac{h_{r}}{B}\sum_{j=1}^{n+k}\sum_{b_{1}=0}^{1}\psi_{0,r,j,b_{1}}e_{j,b_{1}}.$
@@
$\kappa_{1,i,b}=$ $\frac{\eta_{0}}{4Bm^{2}}\left(\sum_{i^{\prime}=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\left(\sum_{r=1}^{2m}\bm{1}\left(\nu_{r,i,b}>0\right)\bm{1}(\nu_{r,i^{\prime},b_{2}}>0)\right)\frac{1}{i^{\prime}}\bar{\alpha}_{i^{\prime},S[i^{\prime}],b_{1},b_{2}}e_{S[i^{\prime}],b_{1}}\right)$
@@
$\left|\frac{\eta_{0}}{8Bmi}\left\langle\left(\sum_{b_{1}=0}^{1}\bar{\alpha}_{i,S[i],b_{1},b}e_{S[i],b_{1}}\right),e_{j,b_{1}}\right\rangle-\frac{\eta_{0}}{8Bmi}\bar{\alpha}_{i,S[i],b_{1},b}\bm{1}(j=S[i])\right|\leq$ $\frac{2\sqrt{\log(200n/\delta)}}{8mn\sqrt{d}}.$
@@
$\left|\langle H_{i,b},e_{j,b_{1}}\rangle-\frac{\eta_{0}}{8Bmi}\bar{\alpha}_{i,S[i],b_{1},b}\bm{1}(j=S[i])\right|\leq\frac{10\eta_{0}}{mn}\frac{\sqrt{nk}}{\sqrt{Bd}}\log(300mnB/\delta).$
@@
$\frac{d\ell(\mathcal{T}(\textbf{b})[i],\textbf{b}[i+1])}{dA}=(-1)^{\textbf{b}[i+1]+1}\sum_{j=1}^{i}\langle\kappa_{1,i,\textbf{b}[i]},e_{j,\textbf{b}[j]}\rangle\frac{\partial\mathrm{softmax}\left(\mathbf{E}(\textbf{b})^{T}A\mathbf{E}(\textbf{b})\right)[j,i]}{\partial A},$
with
@@
$=\sum_{r=1}^{2m}h_{r}W^{(1)}_{r,d+1:2d}\bm{1}\left(\nu_{r,i,\textbf{b}[i]}>0\right)$
@@
$\frac{d\ell(\mathcal{T}(\textbf{b})[i],\textbf{b}[i+1])}{d\mathcal{A}(\mathbf{E}(\textbf{b}))[i]}$ $=(-1)^{\textbf{b}[i+1]}\left(\sum_{r=1}^{2m}h_{r}W^{(1)}_{r,d+1:2d}\bm{1}\left(\nu_{r,i,\textbf{b}[i]}>0\right)\right)$
@@
$=$ $(-1)^{\textbf{b}[i+1]}\sum_{j=1}^{i}\sum_{r=1}^{2m}h_{r}\bm{1}\left(\nu_{r,i,\textbf{b}[i]}>0\right)\langle W^{(1)}_{r,d+1:2d},\mathbf{E}(\textbf{b})[j]\rangle\frac{\partial\mathrm{softmax}\left(\mathbf{E}(\textbf{b})^{T}A\mathbf{E}(\textbf{b})\right)[j,i]}{\partial A}$
@@
$-\frac{dL^{(1)}}{dA}=$ $\frac{1}{B}\sum_{i=n+1}^{n+k}\sum_{j=1}^{i}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i}\alpha_{1,i,j,b_{1},b_{2}}\Delta_{1,i,j,b_{1},b_{2}}e_{j,b_{1}}e_{i,b_{2}}^{T}$
$-\frac{1}{B}\sum_{i=n+1}^{n+k}\sum_{j=1}^{i}\sum_{p=1}^{i}\sum_{b_{1}=0}^{1}\sum_{b_{2}=1}^{1}\sum_{b_{3}=0}^{1}\frac{1}{i^{2}}\mu_{1,j,i,p,b_{1},b_{2},b_{3}}\Delta_{1,i,j,b_{1},b_{2}}e_{p,b_{3}}e_{i,b_{2}}^{T}.$
@@
    $U_{j,p}$ can be bounded in the same way as $T_{j}$ and we have that,
@@
$\left|\mathcal{A}[A^{2}](\mathbf{E}(\textbf{b}))[i]-e_{S[i],\textbf{b}[S[i]]}\right|<1/n^{10}.$
Proof of Lemma 22.
This is a direct combination with Lemmas 21, 32 and 5. ∎
B.4.4 Third Step: Moving MLP in Linear Regime